# DiVerify: Hardening Identity-Based Software Signing with Programmable Diverse-Context Scopes

### v2

@@ 7.2 Mitigating Past Attacks with DiVerify
    ESLint (Type I): The attackers used stolen maintainer credentials to publish a malicious package. With DiVerify, identity verification would have required additional contextual scopes, not just possession of the maintainer’s token. Since the attacker could not produce these scopes, their attempt to propagate the modified package via npm would have failed
  2. 2.
@@
    Midnight Blizzard (Type II): The attackers leveraged OAuth tokens obtained through a compromised identity provider. If the relying services had required DiVerify’s contextual scopes beyond simply accepting valid OAuth tokens, the attacker would have been unable to authenticate through the necessary scope providers, and the attack would have been prevented.
  3. 3.
@@
8 Implementation and Evaluation
In this section, we assess DiVerify ’s practicality by measuring signing/verification times and storage overhead. Given that our implementation is built on existing signing tooling, the only costs for adoption are those of storing and executing new processes and tools within an ecosystem.
8.1 Implementation
Table 2: Implemented Scope Providers(see section˜5.3 for type) Location indicates where provider runs. Size is lines of code via sloc with default parameters.
@@ 8.1 Implementation
Security Key [90] SP Local 28
Table 3: Average Signing & Verification Overhead Introduced by DiVerify (in milliseconds). L1–L3 indicate trust levels determining the signer scopes required during signing.
@@
TUF 1287.25 1162.46 1230.16 1164.88 1217.34 1184.39 1287.41 1274.11 1295.11
We extended the securesystemslib interface for signing and verifying signatures [45] to prototype DiVerify capabilities by making four major changes:
@@
First, we implemented a DiVerify-aware signing to support both signing client and verifier. The client is responsible for retrieving signer scopes, signing and generating remote attestation (section˜5.2.2, and section˜5.2.3); on the verifier side, it enforces scope checks against embedded proofs and package policy (section˜5.2.5). We also implemented separate daemon logic to support operational modes where scope retrieval and signing are delegated to a dedicated background service. We deployed a local RSTUF[72] to securely distribute DiVerify policy(section˜5.4). Second, we adapted the client to support the models described in section˜6. We adapted this client along other identity-based software signing tools such as sigstore-python [82] with similar arguments and parameters to ensure user familiarity.
@@
Third, we utilized existing authentication protocols to provide user scopes. We use the token beacon [11] to request identity tokens from a GitHub OIDC provider. We implemented two static scope providers: a custom one for measuring device fingerprint, and another for attesting[90] the signer’s security key. See Table 2. We provide three bespoke configurations using these scope providers as tighter security settings (L1 to L3).
@@
Finally, we customized the Fulcio certificate authority [81] to incorporate DiVerify proof verification prior to certificate issuance. This required modifying 201 lines of code. Verified proofs are embedded in the issued certificate.
@@
To support remote attestation, we deployed a local Provisioning Certificate Caching Service [15], which supplies quote generation and verification services with Intel SGX collateral: Provisioning Certification Key certificates, Certificate Revocation List, Trusted Computing Base info, and Quoting Enclave identity data. For quote verification, we use Intel’s SGX DCAP Quote Verification tool [16].
Experiment Setup.
@@
We ran our experiments using Docker containers on a 4.0 GHz Intel Xeon Platinum 8580 machine with 499GB of RAM and SGX support, reflecting a realistic industry-grade setup for scalable signing tasks [37]. TEE components were run using Gramine [12], which automatically handles the loading and execution of apps within the enclave.
8.2 Performance Evaluation
@@
We evaluated DiVerify’s performance across the three modes, legacy-compatible, Trusted Authentication, and Core. All experiments used the sample trust level configuration shown in Listing LABEL:lst:diverify-config-structure.
@@
We evaluated different policies to mode mapping with constraints to mitigate the attack vectors in our threat model. To evaluate DiVerify’s performance overhead, we averaged over 10 iterations of artifact signing and verification for each combination of mode/trust level.
8.2.1 Performance Overhead
@@
Table˜3 presents breakdown of time spent across the signing and verification phases. The primary contributor to signing latency is the Fulcio CA. This is evident in Trusted Authentication Mode, where the CA verifies client integrity attestation leading to longer signing times compared to the Legacy, which skips attestation.
@@
Signing using Core offers stronger security guarantees with minimal overhead. It can reduce signing time by up to 13ms compared to the legacy-compatible variant because all operations are handled within the daemon, avoiding round trips to an external CA. Still, both signing modes complete in under 100ms. On the other hand, verification in Core is more expensive, due to additional attestation verification performed by the verifier, resulting to about 11% average latency increase compared to the legacy-compatible mode.
@@
Despite the added security guarantee, DiVerify’s performance remains comparable to state-of-the-art benchmarks [62]. On the verification side, the main cost stems from updating TUF metadata, which as specified, requires multiple round-trips to a remote repository. Core incurs slightly higher overhead than the other modes due to the additional verifier-side effort in validating the client’s integrity attestation. Nevertheless, the overhead remains within acceptable bounds and on par with current signing and verification tools.
8.2.2 Storage Overhead
@@
We measured the size of DiVerify proofs across different operational modes, ranging from 0.4 KB (minimal identity attributes) to 6.7 KB (with full scope and attestation). This size remains fixed for deployments with consistent scope requirements. To understand the practical impact, we analyzed the top 20 most downloaded PyPI packages over the past month [70] and found an average package size to be 1.68 MB. From our sample, DiVerify adds at most 0.39% overhead, which is minimal relative to the package size. While the relative impact decreases for larger packages, the added size remains modest even for smaller ones.
9 Discussion
9.1 DiVerify & Signer Privacy
@@
Our work thus far has focused on security and deployment objectives, not on anonymity or privacy for the signer. DiVerify, as described, ties a lot of identifiable information into the signature bundle, which might concern privacy-conscious signers or organizations. For instance, the certificate will reveal exactly which identity, devices, and scope providers were used. To address this privacy concern, DiVerify can integrate Speranza[52]. Speranza relies on a chain-of-custody proof (i.e., identity co-commitments) from a repository and an IdP. Integrating DiVerify with Speranza would require to adapt scope providers to provide a portion of an identity CoCommitment.
Consequently, while the set of providers remains visible, it becomes impossible to decipher the user who authenticated with them. This effectively provides privacy for signers, ensuring their anonymity when signing.
9.2 Applicability/Deployment Considerations
@@
DiVerify’s design emphasizes backward compatibility (deployment goal (D1)). It builds on standard authentication flows and adds contextual identity “scopes” to signatures, allowing it to plug into existing code signing systems without replacing any infrastructure. Because of this, organizations can introduce DiVerify alongside current workflows; existing verification processes still work (even if they ignore the extra scope data). This means teams can roll out DiVerify incrementally – for example, starting with basic identity checks and gradually adding more stringent scope requirements as they become ready.
@@
Integrating DiVerify requires little changes to existing tooling (satisfying D2). Our demonstration on secure systems library [45] illustrates this. The integration required only 2007 lines of code for the client side signing and verification tooling, and 201 lines of code increase for Fulcio. These modest changes implies maintainers can add DiVerify with very little development effort, keeping the codebase easy to manage.
@@
Because DiVerify leverages existing capabilities, maintainers don’t have to reinvent their environment, just augment it. Another consideration is the choice of scopes for a given context. Our approach is flexible across the claim types identified in 2.2. In automated build systems, DiVerify could be used to ensure that build and signing servers themselves authenticate via scopes (like machine certificates or runner attestations), and that triggers for signing (like a GitHub Actions workflow run) carry a scope proving an authorized user initiated them.
10 Related Work
@@
Role-Based Access Control (RBAC) has long been foundational in managing permissions [77], with widespread adoption in general-purpose systems [57, 56], cloud platforms [41, 2], and social coding environments [21, 22]. In supply chain contexts, Kuppusami et al. [43, 23] propose repository-specific models for permission management, leveraging abstractions like branches and commits. DiVerify differs by mapping permissions across heterogeneous domains without sacrificing security guarantees, particularly in the identity-based software signing context.
Several studies have also identified limitations in RBAC implementations for supply chain security [58, 39], focusing on repository metadata and CI platforms. These insights inform DiVerify’s design, especially in modeling semantics across diverse systems.
@@
Similarly, the application of identity for different contexts has surfaced in general identity work. Perhaps one of the richest lines of work in this regard is that of Verifiable Credentials. Work from [34, 50, 59, 47, 66] introduces the ability to disclose, verifiably, that a property is held by a credential holder. DiVerify provides a similar goal by building upon existing and widely deployed technologies (such as OpenID Connect). Properties such as the ability to commit to a repository or holding ownership of a project can be achieved by off-the-shelf algorithms and tools.
Similarly, applications of distributed identity frameworks (i.e., DiD [89, 5, 88]) can be used in lieu of identity providers. This has been considered in standards such as SCITT as a possible building block for a global software supply chain ledger [29]. Were these standards to become widespread, DiVerify can leverage them. However, it is still necessary to apply the policy logic to merge various providers described in this paper (section˜5.3) to ensure proper capabilities are applied.
@@
Threshold-based authentication schemes, which distribute trust across multiple identity providers [33] using secret sharing [80], have been proposed for enhancing resilience and privacy in federated identity systems. These approaches align with DiVerify’s goal of federated identity verification but typically rely on a trusted dealer. DiVerify avoids this by ensuring no single party can subvert trust, maintaining a decentralized model
11 Conclusion
@@
  • Appendix˜A: An extended version of the assumptions we made.
  • Appendix˜B: Sample DiVerify Artifact Structure
  • Appendix˜C: Sample policy used in our analysis, with signer scope mapped to the expected claims, and the constraint the signature must meet to be trusted.
@@ Appendix A Extended Assumptions
Beyond the attacker models described in section˜3.2, we make standard assumptions from previous literature [7, 62]. First, attackers are able to interpose communications, but standard mechanisms to prevent person-in-the-middle attacks (e.g., TLS) are in place and all parties utilize these. Similarly, attackers are able to replay communication messages to attempt to impersonate roles in the system.
@@ Appendix B Other DiVerify Artifact Samples
Listing 2: Sample Trust Level
@@
Listing 3: Sample DiVerify Proof
@@ Appendix C Test DiVerify Policies
Figure˜5 presents a sample policy used to evaluate whether to trust a package signature.
Figure 5: An example DiVerify policy specifying expected claims. The final field is a Boolean rule constraint under which a package signature is to be trusted. This policy is defined by the software producer and enforced by the verifier, ensuring that only signatures that meet this constraint are accepted, which strengthens the overall trust in the signing process.

### v3

@@ 8.1. Mitigating Past Attacks with DiVerify
    ESLint (Type I): The attackers used stolen maintainer credentials to publish a malicious package. With DiVerify, identity verification would have required additional contextual scopes, not just possession of the maintainer’s token. Since the attacker could not produce these scopes, their attempt to propagate the modified package via npm would have failed.
  2. (2)
@@
    Midnight Blizzard (Type II): The attackers leveraged OAuth tokens obtained through a compromised identity provider. If the relying services had required DiVerify’s contextual scopes beyond simply accepting valid OAuth tokens, the attacker would have been unable to authenticate through the other scope providers, and the attack would have been prevented.
  3. (3)
@@
Existing systems base trust solely on possession of valid credentials or tokens, leaving verifiers blind to compromised identities, providers, or tools. DiVerify prevents these attacks by exposing diverse, verifiable security context to verifiers and enforcing policies over that context, enabling misuse to be detected and rejected at verification time rather than after an attack.
9. Implementation and Evaluation
We now assess DiVerify’s practicality by measuring signing/verification times and storage overhead. Given that our implementation is built on existing signing tooling, the only costs for adoption are those of storing and executing new processes and tools within an ecosystem.
9.1. Implementation
Table 2. Implemented Scope Providers (see section 5.3 for type). Location indicates where provider runs. Size is lines of code via sloc with default parameters.
@@ 9.1. Implementation
Security Key (Yubico, [n.d.]) SP Local 28
Table 3. Average Signing & Verification Overhead Introduced by DiVerify (in milliseconds). L1–L3 indicate trust levels determining the signer scopes required during signing.
@@
We build upon the widely utilized signing provide library Securesystemslib (Lab, [n.d.]) used by tools like TUF(gittuf, 4 29), in-toto (Torres-Arias et al., 2019), and Sigstore (Newman et al., 2022). Securesystemslib provides core identity-based signing and verification abstractions, which allowed us to prototype DiVerify by augmenting existing signing and verification interfaces. We implement DiVerify by making four major changes:
@@
First, we implemented a DiVerify-aware signing to support both signing client and verifier. The client is responsible for retrieving signer scopes, signing and generating remote attestation (section 5.1); on the verifier side, it enforces scope checks against embedded proofs and package policy (section 5.2). We also implemented separate daemon logic to support operational modes where scope retrieval and signing are delegated to a dedicated background service. Second, we adapted the client to support the models described in section 7. We adapted this client along other identity-based software signing tools such as sigstore-python (Sigstore Project, [n.d.]) with similar arguments and parameters to ensure user familiarity.
@@
Third, we utilized existing authentication protocols to provide user scopes. We use the token beacon (Conformance, 0 12) to request identity tokens from a GitHub OIDC provider. We implemented two static scope providers: a custom one for measuring device fingerprint, and another for attesting(Yubico, [n.d.]) the signer’s security key. See Table 2. We advertise supported scopes using configurable trust levels (shown in Listing LABEL:lst:diverify-config-structure) that shows progressively stronger authentication and context signals. These levels range from OIDC-based identity assertions (Level 1), to the inclusion of a hardware-backed security key (Level 2), and to device-bound identity assertions (Level 3).
@@
Finally, we customized Sigstore’s Fulcio certificate authority (Sigstore, [n.d.]) to incorporate DiVerify proof verification prior to certificate issuance. This required modifying 201 lines of code. Verified proofs are embedded in the issued certificate.
@@
To support remote attestation, we deployed a local Provisioning Certificate Caching Service (Corporation, [n.d.]a), which supplies quote generation and verification services with Intel SGX collateral: Provisioning Certification Key certificates, Certificate Revocation List, Trusted Computing Base info, and Quoting Enclave identity data. For quote verification, we use Intel’s SGX DCAP Quote Verification tool (Corporation, [n.d.]b).
Experiment Setup
@@
We ran our experiments using Docker containers on a 4.0 GHz Intel Xeon Platinum 8580 machine with 499GB of RAM and SGX support, reflecting a realistic industry-grade setup for scalable signing tasks (Kalu et al., 2025b). TEE components were built using Gramine (Contributors, [n.d.]), which automatically handles the loading and execution of apps within the enclave.
9.2. Performance Evaluation
@@
We evaluated DiVerify’s performance across the three modes, legacy-compatible, Trusted Authentication, and Core.
@@
We evaluated different policies to mode mapping with constraints to mitigate the attack vectors in our threat model. To evaluate DiVerify’s performance overhead, we averaged over 10 iterations of artifact signing and verification for each combination of the implemented modes and trust levels.
9.2.1. Performance Overhead
@@
Table 3 presents breakdown of time spent across the signing and verification phases. The primary contributor to signing latency is the Fulcio CA. This is evident in Trusted Authentication Mode, where the CA verifies client integrity attestation, leading to longer signing times compared to the Legacy, which skips attestation.
@@
Signing using Core mode offers stronger security guarantees with minimal overhead. It provides less overhead (13ms) compared to the legacy-compatible variant because all operations are handled within the daemon, avoiding round trips to an external CA. Still, both signing modes provide acceptable overhead since they complete in under 100ms. However, verification in Core is more expensive due to additional attestation verification performed by the verifier, resulting to about 11% average latency increase compared to Legacy-Compatible.
@@
Despite the added security guarantees, DiVerify’s performance remains comparable to state-of-the-art benchmarks (Newman et al., 2022). Core mode incurs slightly higher overhead than the other modes due to the additional verifier-side effort in validating the client’s integrity attestation. Nevertheless, the overhead remains within acceptable bounds and on par with current signing and verification tools.
9.2.2. Storage Overhead
@@
We measured the size of DiVerify proofs across different operational modes, ranging from 0.4 KB (minimal identity attributes) to 6.7 KB (with full scope and attestation). This size remains fixed for deployments with consistent scope requirements. To understand the practical impact, we measured the size of the top 20 most downloaded PyPI packages in the past month (PyPI Stats, [n.d.]) and identified an average package size of 1.68 MB. DiVerify adds at most 0.39% overhead to this average size — a minimal cost relative to the package size. While the relative impact decreases for larger packages, the added size remains modest even for the smaller ones.
10. Discussion
In the interest of practicality, there are further considerations and extensions that lie outside of our system as presented thus far.
10.1. DiVerify & Signer Privacy
@@
Our work thus far has focused on security and deployment objectives, not on anonymity or privacy for the signer. DiVerify, as described, ties a lot of identifiable information into the signature bundle, which might concern privacy-conscious signers or organizations. For instance, the certificate will reveal exactly which identity, and scope providers were used. To address this privacy concern, DiVerify can integrate Speranza (Merrill et al., 2023). Speranza relies on a chain-of-custody proof (i.e., identity co-commitments) from a repository and an IdP. Integrating DiVerify with Speranza would require to adapt scope providers to provide a portion of an identity CoCommitment. Consequently, while the set of providers remains visible, it becomes impossible to decipher the user who authenticated with them. This effectively provides privacy for signers, ensuring their anonymity when signing.
10.2. Applicability/Deployment Considerations
@@
DiVerify’s design emphasizes backward compatibility. It builds on standard authentication flows and adds contextual identity “scopes” to signatures, allowing it to plug into existing code signing systems without replacing any infrastructure. Because of this, organizations can introduce DiVerify alongside current workflows; existing verification processes still work (even if they ignore the extra scope data). This means teams can roll out DiVerify incrementally – for example, starting with basic identity checks and gradually adding more stringent scope requirements as they become ready. In other words, current flows are effectively a rather constrained instantiation of DiVerify, as far as existing clients are aware.
@@
Integrating DiVerify requires little changes to existing tooling. Our implementation illustrates this. The integration only required 2007 lines of code for the client side signing and verification tooling, and 201 lines of code increase for Fulcio. These changes imply maintainers can add DiVerify with very little development effort.
@@
Because DiVerify leverages existing software signing architecture, maintainers need not change their environments, but rather augment them. Another consideration is the choice of scopes for a given context. Our approach is flexible across the claim types identified in section 2.2. In automated build systems, DiVerify could be used to ensure that build and signing servers themselves authenticate via scopes (like machine certificates or runner attestations), and that triggers for signing (like a GitHub Actions workflow run) carry a scope proving an authorized user initiated them.
As with multi-factor authentication systems (De Cristofaro et al., 2014), DiVerify does not assume a single notion of authentication effort. Strong security improvements generally require additional, non-substituable evidence, and therefore introduce trade-offs that vary across deployments. DiVerify is designed to make these trade-offs manageable by allowing maintainers to choose which combination of identity scopes are required for acceptance.
11. Related Work
@@
Role-Based Access Control (RBAC) has long been foundational in managing permissions (Sandhu et al., 1996), with widespread adoption in general-purpose systems (Microsoft Azure, 4 21; Microsoft, 4 29), cloud platforms (kubernetes, 4 29; Ansible Project, 2024), and social coding environments (GitHub, 2024; GitLab, 2024). In supply chain contexts, Kuppusami et al. (Kuppusamy et al., 2016; gittuf, 4 29) propose repository-specific models for permission management, leveraging abstractions like branches and commits. DiVerify differs by mapping permissions across heterogeneous domains without sacrificing security guarantees, particularly in the identity-based software signing context.
Several studies have also identified limitations in RBAC implementations for supply chain security (Moore et al., 2023; Koishybayev et al., 2022), focusing on repository metadata and CI platforms. These insights inform DiVerify’s design, especially in modeling semantics across diverse systems.
@@
Similarly, the application of identity for different contexts has surfaced in general identity work. Perhaps one of the richest lines of work in this regard is that of Verifiable Credentials. Prior works show how credential holders can selectively and verifiably disclose properties they possess (Johnson et al., 2023; Lux et al., 2020; Mukta et al., 2020; Lim et al., 2021; OpenID Foundation, nd) DiVerify provides a similar goal by building upon existing and widely deployed technologies (such as OpenID Connect). Properties such as the ability to commit to a repository or holding ownership of a project can be achieved by off-the-shelf algorithms and tools.
Similarly, applications of distributed identity frameworks (i.e., DiD (World Wide Web Consortium, 2022; bcGove, 2023; walt.id, 2023)) can be used in lieu of identity providers. This has been considered in standards such as SCITT as a possible building block for a global software supply chain ledger (IETF, 2022). Were these standards to become widespread, DiVerify can leverage them. However, it is necessary to apply the policy logic to merge various providers described in this paper (section 5.3) to ensure proper capabilities are applied.
@@
Threshold-based authentication schemes, which distribute trust across multiple identity providers (Ito et al., 2017) using secret sharing (Shamir, 1979), have been proposed for enhancing resilience and privacy in federated identity systems. These approaches align with DiVerify’s goal of federated identity verification but typically rely on a trusted dealer. DiVerify avoids this by ensuring no single party can subvert trust, maintaining a decentralized model
12. Conclusion
@@
  • Appendix A: An extended version of the assumptions we made.
  • Appendix B: Sample DiVerify Artifact Structure
  • Appendix C: Sample policy used in our analysis, with signer scope mapped to the expected claims, and the constraint the signature must meet to be trusted.
@@ Appendix A Extended Assumptions
Beyond the attacker models described in section 3.2, we make standard assumptions from previous literature (Cappos et al., 2008; Newman et al., 2022). First, attackers are able to interpose communications, but standard mechanisms to prevent person-in-the-middle attacks (e.g., TLS) are in place and all parties utilize these. Similarly, attackers are able to replay communication messages to attempt to impersonate roles in the system.
@@ Appendix B Other DiVerify Artifact Samples
Listing 1: Sample Trust Level
@@
Listing 2: Sample DiVerify Proof
@@ Appendix C Test DiVerify Policies
Figure 5 presents a sample policy used to evaluate whether to trust a package signature.
Figure 5. An example DiVerify policy specifying expected claims. The final field is a Boolean rule constraint under which a package signature is to be trusted. This policy is defined by the software producer and enforced by the verifier, ensuring that only signatures that meet this constraint are accepted, which strengthens the overall trust in the signing process.